The rate constant for the reaction 2NO → O2 + N2 is 8.0×10−3 Atm−1min−1 at 800...

  1. The rate constant for the reaction 2NO → O2 + N2 is 8.0×10−3 Atm−1min−1 at 800 K. How long will it take the concentration of NO to drop from 0.85 Atm to 0.35 Atm?

    1. a) 110 min

    2. b) 250 min

    3. c) 180 min

    4. d) 210 min

    5. e) 520 min

  2. For the second order reaction 2SiH4 → Si2H6 + H2

    has a rate constant of 5.5×10−4 Atm−1 s−1 at 230oC. If a vessel contains 1.326 Atm of SiH4 after a reaction time of 1200 s, what was the initial SiH4 partial pressure?

    1. a) 10.6 Atm

    2. b) 0.707 Atm

    3. c) 2.565 Atm

    4. d) 7.29 Atm

    5. e) 87.4 Atm

    6. e) 990 min
Answer #1

1)
we have:
[NO]o = 0.85 atm
[NO] = 0.35 atm
k = 8*10^-3 atm-1.min-1

Unit of k suggests that it is 2nd order reaction.

use integrated rate law for 2nd order reaction
1/[NO] = 1/[NO]o + k*t
1/(0.35) = 1/(0.85) + 8*10^-3*t
2.857 = 1.176 +8*10^-3*t
8*10^-3*t = 1.681
t = 210 min
Answer: d

Only 1 question at a time please